Shift summary for Eve, Dec. 28, 1999

Angela Bellavance (bellavan@rice.edu)
Wed, 29 Dec 1999 00:01:44 -0600 (CST)

Shift Summary for eve shift, 28/Dec/99

On Shift: V. Jejer, A. Bellavance

Hours of beam delivered at all intensities: 8.0
Hours of beam delivered at nominal intensity: 8.0
Hours of E799 data (special runs not included): 7.7
Hours of special runs (laser scans, muon runs...): 0.2

Accounting of lost time: NM3 interlock trip, changing modes, TeV aborts

========================================================================
Runs ended this shift:
Run 15313 - BEAM_MUON
Run 15314 - ongoing - BEAM_799

Tapes brought to FCC this shift: UPA274-6
UPB274-6
UPC274-6
UPS092
UPH049-50
UPF067
UPK035
UPT022

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E799 datataking after a muon run. Semi-frequent scaler alarms
throughout. Occasional HCC inefficient alarms.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15312 15:30 Run just ends at shift switchover with full tapes.
We set up to do a magnet flip and muon run.
15:36 Mounted UPF068.
15313 1 15:41 Start muon run.
15313 15:51 Stop muon run. Request MCR to go back to data
taking. Message on Ch. 13: "Neutrino primary
permit dropped. Investigating."
16:06 Ch.13: "Beam to users", but we are not in data
taking mode yet.
15314 1 16:08 In correct mode. Start run.
17:48 Atev3 dies. Reset.
18:03 Mounted UPM051.
89 18:07 Ktevana errors. One spill missed while KFRENDing.
90 18:08 1/2 spill abort. Fine next spill.
169 19:55 TeV abort.
175 20:03 SY energy conservation.
20:05 Mounted UPK036.
185 20:19 SY energy conservation.
191 20:24 Ktevana errors, KFREND.
193 TeV abort.
21:00 TeV abort.
21:22 TeV abort.
243 21:34 Ktevana errors, KFREND.
21:58 SY energy conservation.
262-4 22:00 Three spills w/ messy spill structures.
275 22:17 TeV abort.
279 22:21 TeV abort.
290 22:37 Warnings from Val about a ped. problem and laser
illumination for channel 1081. BobT paged.
292 22:40 TeV abort.
297 22:45 BobT suggests bad laser ill. may be pipeline
problem. Do a pipeline reset and KFREND. Channel
1081 returns.
302 Channel 1081 gain drifts by 50% but is ok again by
next spill.
23:10 Notice a high-pitched whine. Looks like a fan in
rack 161 (second row from top) is dying. This
will need to be looked at next down time, or when
it dies, whichever comes first.
318 23:15 SY energy conservation.
323 23:22 TeV abort.
347 23:54 TeV abort.

========================================================================
Follow up Notes on Previous Shift

-muon run done, both AN and SR magnets flipped to NEG.

========================================================================
Items to Follow up on Next Shift

-keep an eye (ear?) on the fan in rack 161.
-SDAQ needs to be fresh-started before the next run.
-page Sasha Golossanov if there is a 1+ hour shutdown (not owl shift).

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke